2. The slant height of a frustum of a cone is 4 cm and the perimeters (circumference) of its circular ends are 18 cm and 6 cm. Find the surface area of the frustum.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Given, Slant height (l) = 4 cm Circumference of upper circular end of the frustum = 18 cm ∴ 2πr1 = 18 Or, r1 = 9/π Similarly, circumference of lower end of the frustum = 6 cm ∴ 2πr2 = 6 Or, r2 = 6/π Now, CSA of frustum = π(r1+r2) × l = π(9/π+6/π) × 4 = 12×4 = 48 cm2